OpenStudy (anonymous):

solve the system of equations by substitution 8y-1=x and 3x-2y

OpenStudy (anonymous):

to solve by substitutions, we must solve one of the equations for a single variable, and use that answer to substitute in for the other equations for example x + y = 4 2x - y = 0 (these are made up by the way) we would solve the first equation x+y = 4 for a single variable so subtract y from both sides x = 4-y so now we would plug (4-y) wherever there an x in the second equation 2(4-y) - y = 0 8 - 2y - y = 0 8 - 3y = 0 now solve for y -3y = -8 y = 8/3 so now that we know what y is, plug your value for y in the equation that we solved for x originally x = 4 - y x = 4 - (8/3) x = 4/3 so x = 4/3 and y = 8/3 apply the same concept to your problem. just follow it step by step